OpenStudy (anonymous):

Simplify the following expression. 7x - 4y + 5z - 2x - 10y + z

jimthompson5910 (jim_thompson5910):

7x - 2x is ???

OpenStudy (anonymous):

5x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

-4y - 10y is ???

OpenStudy (anonymous):

-14

jimthompson5910 (jim_thompson5910):

-14y, but you're on the right track

jimthompson5910 (jim_thompson5910):

finally, 5z + z is ???

OpenStudy (anonymous):

6z?

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

so 7x - 4y + 5z - 2x - 10y + z simplifies to 5x - 14y + 6z

OpenStudy (anonymous):

Combine like terms. 7x-2x = 5x -4y -10y= -14y 5z+z= 6z 5x-14y+6x should be the answer. All you have to do is combine like terms as in the ones with the same variable.

OpenStudy (anonymous):

Didn't want to enter it in the beginning because you were explaining it :P @jim_thompson5910

jimthompson5910 (jim_thompson5910):

thanks for being considerate, i appreciate it AmauryP

OpenStudy (anonymous):

No problem(: